### v1

# HOLISMOKES XVI: Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal

e-mail: stefan.schuldt@unimi.it 22institutetext: INAF – IASF Milano, via A. Corti 12, I-20133 Milano, Italy 33institutetext: Aix-Marseille Université, CNRS, CNES, LAM, Marseille, France 44institutetext: Purple Mountain Observatory, No. 10 Yuanhua Road, Nanjing, Jiangsu, 210033, People’s Republic of China 55institutetext: Technical University of Munich, TUM School of Natural Sciences, Physics Department, James-Franck-Straße 1, 85748 Garching, Germany 66institutetext: Max-Planck-Institut für Astrophysik, Karl-Schwarzschild Straße 1, 85748 Garching, Germany
HOLISMOKES XVI: Lens search in HSC-PDR3 with a neural network committee and post-processing for false-positive removal
We have carried out a systematic search for galaxy-scale lenses exploiting multi-band imaging data from the third public data release of the Hyper Suprime-Cam (HSC) survey with the focus on false-positive removal, after applying deep learning classifiers to all 110 million sources with -Kron radius above 08. To improve the performance, we tested the combination of multiple networks from our previous lens search projects and found the best performance by averaging the scores from five of our networks. Although this ensemble network leads already to a false-positive rate (FPR) of 0.01% at a true-positive rate (TPR) of 75% on known real lenses, we have elaborated techniques to further clean the network candidate list before visual inspection. In detail, we tested the rejection using SExtractor and the modeling network from HOLISMOKES IX, which resulted together in a candidate rejection of 29% without lowering the TPR. After the initial visual inspection stage to remove obvious non-lenses, 3,408 lens candidates of the 110 million parent sample remained. We carried out a comprehensive multi-stage visual inspection involving eight individuals and identified finally 95 grade A (average grade ) and 503 grade B () lens candidates, including 92 discoveries showing clear lensing features that are reported for the first time. This inspection also incorporated a novel environmental characterization using histograms of photometric redshifts. We publicly release the average grades, mass model predictions, and environment characterization of all visually inspected candidates, while including references for previously discovered systems, which makes this catalog one of the largest compilation of known lenses. The results demonstrate that (1) the combination of multiple networks enhances the selection performance and (2) both automated masking tools as well as modeling networks, which can be easily applied to hundreds of thousands of network candidates expected in the near future of wide-field imaging surveys, help reduce the number of false positives that is the main limitation in lens search to date.
Key Words.:
gravitational lensing: strong methods: data analysis catalogs1 Introduction
Strong gravitational lensing emerged in the past decades as a powerful tool to probe galaxy evolution and cosmology. It allows us to obtain in a very precise way the total mass (i.e., baryonic and dark matter) of the galaxy or galaxy cluster acting as the lens (e.g., Bolton et al., 2008; Shu et al., 2017; Caminha et al., 2019) and, by assuming that mass follows light, we can disentangle the mass components and obtain unique insights into the dark matter (DM) distribution (e.g., Schuldt et al., 2019; Shajib et al., 2021; Wang et al., 2022) or DM substructure (e.g., Enzi et al., 2024; Ertl et al., 2024; Lange et al., 2024). Thanks to the lensing magnification, strong lensing also allows us to study high-redshift sources not visible otherwise (e.g., Shu et al., 2018; Vanzella et al., 2021; Meštrić et al., 2022; Stiavelli et al., 2023; Morishita et al., 2024).
In case of a time variable background object, such as a supernova (SN) or a quasar, time delays can be measured (e.g., Courbin et al., 2018; Millon et al., 2020) and exploited for competitive measurements of the value of the Hubble constant (Refsdal, 1964). Given the rarity of SNe and strong lensing events, this time-delay cosmography (TDC) technique was mostly carried out with quasars (e.g., Wong et al., 2020; Acebron et al., 2022; Shajib et al., 2022; Acebron et al., 2024). To date only three strongly lensed SNe are known with time delays usable for a precise measurement of : SN Refsdal lensed by the cluster MACS J1149.52223 (e.g., Grillo et al., 2018; Kelly et al., 2023; Grillo et al., 2024), SN H0pe strongly lensed by the cluster PLCK G165.767.0 (Pascale et al., 2025), and SN Encore with SN Requiem both lensed by the same cluster MACS J01382155 (Ertl et al. in prep., Suyu et al. in prep., Pierel et al. in prep). To prepare a systematic search with the current and upcoming wide-field imaging surveys, we initiated the Highly Optimized Lensing Investigations of Supernovae, Microlensing Objects, and Kinematics of Ellipticals and Spirals (HOLISMOKES Suyu et al., 2020) program. As a precursor for the Legacy Survey of Space and Time (LSST) of the Vera C. Rubin Observatory (Ivezic et al., 2008; Ivezić et al., 2019), we are currently exploiting data from the Hyper Suprime-Cam (HSC), which are expected to be very similar.
In Cañameras et al. (2021, hereafter C21) we presented a residual neural network to search broadly for any static lens, and complemented this in Shu et al. (2022, hereafter S22) with lenses/deflectors at relatively high redshift () to better cover the whole redshift range. Both projects relied on a single convolutional neural network (CNN) and targeted HSC images of the second public data release (PDR2). In this work, we combine multiple networks into an ensemble network, relax our restrictions on filter coverage to increase the observed sky area (requesting bands instead of ), and consider data from PDR3 with slightly deeper and larger footprint.
While the resulting sample of network candidates from HSC is small enough for visual inspection, this will not be the case for LSST, Euclid (Laureijs et al., 2011), Roman (Spergel et al., 2015) and the Chinese Space Station Telescope (Gong et al., 2019), delivering more than a billion images, even with a citizen science approach (e.g., Holloway et al., 2024, Holloway et al in prep., Walmsley et al. in prep.). Consequently, we urgently need further automated ways to lower the false-positive rate (FPR) before an unavoidable visual inspection.
In this paper, we show our deep learning ensemble network and, following C21, apply it to cutouts of any object with -Kron radius above 08 that pass standardised HSC image quality flags (see C21 for details). We explore two different approaches to reject false positives: (1) we run SExtractor to reject images with artefacts or without a real astrophysical source in the cutout, followed by a exclusion through HSC-pixel flags, and (2), we run the residual neural network from Schuldt et al. (2023a, hereafter S23a) to reject systems based on the mass model predictions. Specifically, Sect. 2 presents the tested network committee and their performances on known real lenses. We then describe the two approaches of contaminant rejection in Sect. 3, before we carry out a visual inspection as described in Sect. 4 and present the newly discovered lens candidates. Finally, we give a summary and conclusion in Sect. 5. We exploit trained networks from C21, S22, and Cañameras et al. (2024, hereafter C24), and adopt consequently also a flat CDM cosmology with = 0.32 (Planck Collaboration et al., 2020) and (Bonvin et al., 2017).
2 Inference with network committees
We followed the approach described in C21 to classify the 110 million galaxies from HSC Wide PDR3 with -band Kron radius 08 and optimize the contamination rate of the candidate strong lens sample. Based on test sets drawn from HSC Wide PDR2 images and designed to closely match a real classification setup, Cañameras et al. (2024) show that the purity and overall classification performance are significantly improved with committees of multiple neural networks (see also e.g., Andika et al., 2023). The highest gain is obtained when combining networks trained on different ground-truth data sets, with different prescriptions for the parameters’ distributions over the mock lenses used as positive examples. Taking the average or multiplication of output scores from networks having little overlap in false positives due to their internal representations is best for improving the true-positive rate (TPR) at low FPR.
We investigated several combinations of neural networks chosen among the best performing ResNet and classical CNNs from C21, S22, and C24. After classifying the 110 million HSC PDR3 image cutouts with these networks, we compared the TPR over three independent sets of confirmed and candidate strong lenses as a function of the total number of lens candidates predicted by the committee. This allowed us to find the optimal committee that maximizes the TPR at a given output sample size (i.e. corresponding to a fixed human inspection time). The first set of test lenses (set 1) includes the 1,249 galaxy-scale grade A or B candidates from the SuGOHI papers 111see https://www-utap.phys.s.u-tokyo.ac.jp/ oguri/sugohi/. The second set (set 2) focuses on 201 galaxy-scale grade A from SuGOHI, which are also included in set 1. The third set (set 3) corresponds to the cleaned set of 178 galaxy-scale grade A or B lens candidates used in C21 C24, which excludes visually identified systems with multiple lens galaxies, or significant perturbation from the environment.
We investigated various combinations of different networks by taking the average or the product of their individual network scores. As examples, the performance of combinations with two, three, four, and five networks are shown in Fig. 1. The best committee that maximizes the TPR in all three test sets at fixed number of candidates was obtained by averaging the scores of five individual networks: the ResNet from C21, Classifier-1 from S22, and three additional ResNets from C24, as illustrated in Fig. 2. We adopted a threshold on the average score of 0.55, which corresponds to 22,393 strong-lens candidates, and a TPR of 51%, 54%, and 75% over test set 1, 2, and 3, respectively. Since the TPR curves reach a plateau for all of the three considered test sets, decreasing the score threshold to include 50,000 additional candidates would have improved the TPR by only 4–5%, which does not justify the substantial increase in human inspection time. Having various test sets was important to check that the plateau in TPR is reached irrespective of the exact set-up of the test set. Since set 3 is representing best the galaxy-galaxy scale lenses we target with this network committee (e.g., set 3 is cleaned from group and cluster lenses), the network committee shows the best performance on this set as expected. Finding about 20,000 candidates among 110 million sources from the parent sample is consistent with the FPR in the range 0.01–0.03% predicted in C24 for networks committees, at a TPR of 75% evaluated among set 3.
The five networks are all from the same type of architecture and include residual neural network blocks (He et al., 2015). Specifically, the Classifier-1 from S22 is based on the CMU DeepLens package from Lanusse et al. (2018), while the other four networks are adapted from the ResNet-18 architecture. The networks were however trained on different ground-truth data sets. We introduce the general properties of the training sets and refer the reader to C21, S22, and C24 for further details.
For all five networks, the realistic mock lenses used as positive examples were simulated with the pipeline described in Schuldt et al. (2021, 2023b). Briefly, the pipeline paints lensed arcs on HSC images of massive luminous red galaxies (LRGs) using Singular Isothermal Ellipsoid (SIE) lens mass profiles, and SIE parameter values inferred from SDSS spectroscopic redshifts and velocity-dispersion measurements. After the high-redshift background sources are drawn from the Hubble Ultra-Deep Field (HUDF, Inami et al., 2017), mock lensed arcs are computed with Glee (Suyu & Halkola, 2010; Suyu et al., 2012b), and convolved with the HSC point spread function, before coaddition with the lens galaxy cutout. Positive examples used to train the ResNet from C21 and Classifier-1 from S22 were drawn from the same parent set of mocks, with (i) nearly-uniform Einstein radius distribution between 075 and 25, (ii) a boosted fraction of lens galaxies at with respect to the parent SDSS sample, (iii) lensed images having , and (iv) a minimal ratio between the brightest arc pixel and the local sky background over the lens LRG cutout. The other three high-performing networks from C24 that are part of the committee were trained on mocks produced with a similar procedure, but with (i) a boosted fraction of red HUDF sources, (ii) a natural Einstein radius distribution between 075 and 25 instead of uniform distribution, and (iii) no lower limit on and a balanced fraction of double and quad configurations. This corresponds to sets labeled L1, L4, and L6 in C24.
In terms of negative examples, four out of the five networks include a mix of 33% spirals, 27% isolated LRGs without arcs, 6% groups, and 33% random galaxies over the HSC footprint, as defined for set N1 in C24. The fifth network, namely Classifier-1 from S22, was primarily targeting high-redshift strong-lenses and trained on negative examples drawn from a parent sample with red () and () colors. All five networks were trained and validated on images in bands, such that we require only the availability of bands in HSC while the image in or can be missing.
3 Cleaning the output candidate list
Before conducting a visual inspection (see Sect. 4) to identify the high-quality strong lens candidates, the catalog of 22,393 candidates was post-processed to remove obvious artefacts and non-lenses. Here we applied mainly two criteria, as detailed below. An overview table of the stages and resulting sample sizes is given in Table 1.
Description | Sample size | Section |
Parent sample (in million) | 110 | 2 |
Network candidates | 22,393 | 2 |
After SExtractor cleaning | 19,820 | 3.1 |
After pixel level cuts | 18,712 | 3.1 |
After CNN lens model cleaning | 15,919 | 3.2 |
After excluding duplicates () | 14,448 | 4 |
After excluding previous inspected systems | 11,773 | 4 |
Inspected in binary classification round | 11,874 | 4.1 |
Inspected by first team | 3,408 | 4.2 |
Recentered after first team’s inspection | 160 | 4.2 |
Inspected by second team | 686 | 4.2 |
Total inspected systems listed in Table 2 | 14,152 | 4.4 |
New published grade A (B) lens candidates | 9 (83) | 4.4 |
New inspected grade A (B) lens candidates | 42 (187) | 4.4 |
All grade A (B) lens candidates | 95 (503) | 4.4 |
Grade A (B) lens candidates with OD | 22 (66) | 4.5 |
Grade A (B) lens candidates with OD | 18 (70) | 4.5 |
New inspected grade A with good model | 28 (67%) | 4.5 |
New inspected grade B with good model | 164 (88%) | 4.5 |
Inspected grade A or B with good model | 192 (84%) | 4.5 |
3.1 Cleaning using SExtractor and HSC pixel level flags
A substantial fraction of contaminants correspond to cutouts with residual background emission in one or all bands (see some examples in the top row of Fig. 3). These cutouts were identified using source masking and estimates of the sky background with SExtractor (Bertin & Arnouts, 1996), leaving 19,820 candidates without loss of any test lens. This step was then refined with information from the pixel-level flags inferred by the HSC pipeline. We searched for optimal cuts to remove contaminants using the flags in bands over the 72 72 pixel images. First, we discarded cutouts having 90% pixels with flags 512 and 1024, corresponding to pixels located near a bright object. This securely excluded cutouts affected by bright neighboring stars within a few tens of arcseconds. Secondly, we removed cutouts with 99% pixels flagged as “detected pixels”, which are more likely to correspond to crowded fields such as stellar clusters, or star-forming clumps within extended disks, than strong gravitational lenses. These two criteria decreased the sample to 18,712 candidates.
3.2 Cleaning using modeling CNN
In parallel, we applied the ResNet from S23a to all 22,393 network candidates. This modeling network is trained on realistic mock images created in a similar way as those used for training the committee network, and obtained great performance in measuring the Einstein radii of lens systems when compared to gleeauto.py (Schuldt et al., 2023b), a code that relies on the well-tested lens modeling software Glee (Suyu & Halkola, 2010; Suyu et al., 2012a). The modeling network predicts the lens mass center and , the lens mass ellipticity and , and Einstein radius of a SIE profile as well as an external shear ( and ), and the corresponding 1 uncertainties.
Since the modeling network is trained on solely lens images (i.e. not in combination with lens classification as in Andika et al. (2024)), it is forced to provide reasonable model parameters (e.g., [05,) even if the given image is clearly not a lens. However, for such a given non-lens image, the modeling network is uncertain and the predicted errors can be significantly higher. Consequently, we can use this fact to reject false-positives from the classification network. To be conservative, we define the cuts here such that we do not exclude any lens candidate from our test set as shown in Fig. 4, which results in an additional rejection of 2,794 systems. We show some examples of rejected candidates in the bottom row of Fig. 3.
In sum, we reject around 29% of the network candidates as false positives by keeping the same TPR on our test sample. This cleaning is performed through automated and fast scripts and thus are also scalable for sample sizes expected from ongoing and upcoming wide-field imaging surveys which are two orders of magnitude higher. While excluding of the contaminants by keeping the same TPR is already a good improvement in the purity, we note that for significantly higher samples more stringent cuts can be applied to reject more contaminants with the downside of possibly excluding some lenses.
4 Visual inspection
Before the visual inspection of the remaining network candidates from Sect. 3, we excluded duplicates among the network candidates. Although CNNs are known to be translation invariant, we only excluded duplicates within two pixels when creating the 110 million parent sample to be very conservative. This ensures that we will not miss any candidate from the network due to a lower score if the lens is not well centered or partly out of the cutout. However, for the visual inspection, we removed duplicate cutouts within 2″. In order to preserve lines of sight with the most extended, and most likely deflector galaxy in the center, we rank ordered the catalog entries by decreasing -Kron radius and removed duplicates of lower rank. Excluding these duplicates reduces the human inspection time further, while the moderate cut of 2″ensures that we will not shift any multiple image out of the inspected cutout that would risk missing the identification.
Furthermore, we excluded additional image stamps previously inspected as part of our HOLISMOKES strong lens searches in earlier HSC data releases (C21; S22; Schuldt et al. (2025), hereafter S25) to lower the need in human resources, as we grade the systems in a similar way and would therefore have similar resulting grades. We only kept 50 lenses (high-quality grade A or B lens candidates, which we simply refer to as “lenses” hereafter) and 50 non-lenses from C21 and S25 relying both on the same network, as well as 50 lenses and 50 non-lenses from S22 for comparison222we use here a lower limit of the threshold score to include also a few systems that would be usually excluded with the cut at .. This resulted in a catalog of 11,874 candidates for visual inspection.
The visual inspection follows closely the procedure proposed by S25 and was conducted jointly with candidates presented by Andika et al. (2024). In short, we first carried out a calibration round using 200 system inspected by all eight inspectors333The visual inspectors are in alphabetical order by last name: I. T. A., S. B., R. C., C. G., A. M., S. S., S. H. S., and S. T. since we introduced several new features in the grading tool as detailed below. As previously, we use a three-panel image to show two different scalings of the bands, and, in contrast to previous inspections, an image of the bands to ease the identification of lensed quasars. While these image stamps have a size of , we additionally show stamps in the same filters and scaling as in the smaller cutout.
4.1 Binary inspection
We then split the catalog among ourselves, such that each system is inspected by two graders in a binary classification to exclude rapidly the majority of obvious non-lenses. As mentioned above, the catalog contains 200 systems from our previous searches, with 100 lenses and 100 non-lenses. Of the 100 lenses (or lens candidates), we recovered 98/100 with previous grade A or B and only missed two systems, namely, systems HSC J222002+060506444Listed as HSC J2220+0605 in C21. and HSC J090929+010030 which previously obtained both an average grade of 1.6 while the lower limit for grade B is 1.5. Of the 100 non-lenses, we forwarded 59/100, which, however, is understandable as several of these galaxies obtained previously an average grade slightly below 1.5 and we aimed on forwarded now all systems above a grade of 1 in the final grading scheme. Interestingly, we also forwarded 12/100 with previously average grade 0. Since this first stage is graded conservatively to rule out obvious non-lenses and there is subsequently another round of inspection for refinement, this result on the 200 systems is overall very good.
4.2 Multi-class inspection
In the next round, four individuals inspected the remaining 3,408 systems from the binary classification. At this stage, each inspector provided one of four possible grades ( corresponding to “no lens”, to “possible lens”, to “probable lens”, and to “definite lens”) and voted in case the lens is significantly offset from the cutout center. In total, 309 systems obtained the “off-center” flag by at least one person, such that the lens center was subsequently corrected. A re-centering at this stage was crucial for the subsequent round of grading by the other four remaining graders, as we also showed in the central image the lens center and Einstein radius predicted by the ResNet from S23a (see bottom row of Fig. 3 for examples).
Based on the average grade obtained from these four grades, we forwarded all systems with either or but standard deviation of the four grades above 0.75. Including those with lower grade but high discrepancy increases the forwarded sample significantly, but minimises the possibility of missing any good lens candidate. This resulted in 686 candidates, including 160 systems that got shifted, which were inspected by the remaining four graders. In addition to the grading and “off-center” flag from the previous round, these four individuals were tasked to also indicate if the predicted lens center or Einstein radius is significantly mis-predicted. Furthermore, for each candidate, two histograms of photometric redshifts from the lens candidate environment, one up to and the other up to , were shown in order to obtain a classification of the environment as well. As examples, the histograms of one system in an overdensity (bottom row) and one in the field (top row), are shown in Fig. 5. This allows us to assess if the predicted mass model, which we provide in Table 2 (see also Table 3 for explanations), is reliable and to characterise their environment, while increasing the visual inspection time only very little.
Name | RA [deg] | Dec [deg] | [″] | Model | References | ||||||
(1) | (2) | (3) | (4) | (5) | (6) | (7) | (24) | (25) | (30) | (31) | (37) |
HSCJ10040031 | 8 | 1.44 | 0.15 | Y | 1.05 | A23 | |||||
HSCJ23050002 | 8 | 1.32 | 0.12 | Y | 0.49 | W18 C21 | |||||
HSCJ22420011 | 8 | 1.51 | 0.11 | Y | 0.38 | S18 R22 | |||||
HSCJ12360035 | 8 | 0.90 | 0.06 | Y | 0.49 | S22 | |||||
HSCJ01020159 | 8 | 1.38 | 0.09 | N | 0.95 | J19 C21 | |||||
Column | Header | Description |
(1) | Name | name for grade A and B lens candidates |
(2) | RA [deg] | Right Ascension in degrees of the inspected candidate |
(3) | Dec [deg] | Declination in degrees of the inspected candidate |
(4) | p | average score predicted by the network committee |
(5) | G | average grade between 0 and 3 obtained though visual inspection (see Sect. 4) |
(6) | sigma_G | standard deviation of obtained visual inspection grades. |
(7) | N_graders | number of visual inspectors (see Sect. 4) |
(8) | _kronflux_radius | Kron radius in the band provided by HSC used for source selection (see Sect. 2) |
(9) | Binary | Flag to indicate sources inspected only in the binary stage (two different graders) |
(10) | Round1 | Flag to indicate systems inspected also in the first round (four different graders) |
(11) | Round2 | Flag to indicate systems inspected also in the second round (eight different graders) |
(12) | A24 | Flag to indicate if system jointly inspected with candidates discovered by Andika et al. (2024) |
(13) | S25 | Flag to indicate systems discovered by this network committee but not re-inspected. Instead, we report the |
visual inspection grade from our earlier work S25 | ||
(14) | S22 | Flag to indicate systems discovered by this network committee but not re-inspected. Instead, we report the |
visual inspection grade from our earlier work Shu et al. (2022) | ||
(15) | C21 | Flag to indicate systems discovered by this network committee but not re-inspected. Instead, we report the |
visual inspection grade from our earlier work Cañameras et al. (2021) | ||
(16) | x_med | -center coordinate predicted by the modeling network from Schuldt et al. (2023a, hereafter S23a) |
(17) | x_err | 1- value for the -center coordinate predicted by the modeling network from S23a |
(18) | y_med | -center coordinate predicted by the modeling network from S23a |
(19) | y_err | 1- value for the -center coordinate predicted by the modeling network from S23a |
(20) | ex_med | -component of the complex ellipticity predicted by the modeling network from S23a |
(21) | ex_err | 1- value for the -component of the complex ellipticity predicted by the modeling network from S23a |
(22) | ey_med | -component of the complex ellipticity predicted by the modeling network from S23a |
(23) | ey_err | 1- value for the -component of the complex ellipticity predicted by the modeling network from S23a |
(24) | rE_med | Einstein radius value of the given candidate predicted by the modeling network from S23a |
(25) | rE_err | 1- value of the Einstein radius predicted by the modeling network from S23a |
(26) | gam1_med | -component of the external shear predicted by the modeling network from S23a |
(27) | gma1_err | 1- value of the -component predicted by the modeling network from S23a |
(28) | gam2_med | -component of the external shear predicted by the modeling network from S23a |
(29) | gam2_err | 1- value of the -component predicted by the modeling network from S23a |
(30) | Model | Flag if the model predict is reliable, only for systems inspected in the second round |
(31) | z | photometric redshift value from the catalog compiled catalog by S25 |
based on DEmP (Hsieh & Yee, 2014), Mizuki (Tanaka et al., 2018), and NetZ (Schuldt et al., 2021) | ||
(32) | OD_vis | Flag if the system falls into a significant overdense region |
(33) | OD_z | Flag if the system is in a significant overdense environment according to the criteria defined by S25 |
(34) | N_max | Peak of the photo- histogram following the procedure of S25 |
(35) | zlow | Lower bound of N_max in the photo- histogram indicating the redshift of the overdensity |
(36) | Ntot | Sum of objects with photo- within a box of 200″on a side. |
(37) | References | List of publications that report the inspected candidate (within 5″) as lens candidate according to the |
HOLISMOKES Suyu et al. (2020) lens compilation with status of the publication, using B04 for | ||
Bolton et al. (2004), C07 for Cabanac et al. (2007), B08 for Bolton et al. (2008), G14 for Gavazzi et al. (2014), | ||
H15 for Holwerda et al. (2015), M16 for More et al. (2016), P16 for Paraficz et al. (2016), S16 for | ||
Shu et al. (2016), D17 for Diehl et al. (2017), J17 for Jacobs et al. (2017), S18 for Sonnenfeld et al. (2018), | ||
W18 for Wong et al. (2018), J19 for Jacobs et al. (2019), L19 for Li et al. (2019), P19 for Petrillo et al. (2019), | ||
H20 for Huang et al. (2020), C20 for Chan et al. (2020), Ca20 for Cañameras et al. (2020), Cao20 for | ||
Cao et al. (2020), L20 for Li et al. (2020), J20 for Jaelani et al. (2020), S20 for Sonnenfeld et al. (2020), C21 | ||
for Cañameras et al. (2021), H21 for Huang et al. (2021), L21 for Li et al. (2021), T21 for Talbot et al. (2021), | ||
R22 for Rojas et al. (2022), S22 for Shu et al. (2022), Sa22 for Savary et al. (2022), St22 Stein et al. (2022), | ||
W22 for Wong et al. (2022), Z22 for Zhong et al. (2022), A23 for Andika et al. (2023), J24 for | ||
Jaelani et al. (2024), G24 for Grespan et al. (2024), St24 for Storfer et al. (2024), S25 for Schuldt et al. (2025), | ||
ML for the master lens catalog at http://admin.masterlens.org, and ‘Guoyou Sun’ corresponds to | ||
candidates identified by an amateur astronomer, Guoyou Sun, through visual inspections of HSC cutouts (see | ||
http://sunguoyou.lamost.org/glc.html.). |
4.3 Comparison to previous inspected candidates
As mentioned above, we included 200 systems that we inspected already in previous works. From the 100 candidates that we previously excluded (i.e. that obtained an average grade ), only one system obtained now an average grade above 1.5 and is consequently listed as grade B candidate. From the sample of previously classified grade A and B lenses, we notice a tendency towards lower grades consistently among most graders. While there were some changes compared to previous inspections, such as moving from PDR2 to PDR3 and showing two color images and one color image instead of three color images, we ruled these aspects out by directly comparing some inspected images from S25 and this work. We note also that the applied scalings depend on the pixel values of the given cutout, such that a shifted image appears slightly different. Additionally, in S22 we focused on high- lenses and applied slightly different scaling functions, but find this to be not a major reason. Furthermore, we speculate that this tendency may come from the relatively pure sample (i.e., a higher proportion of lenses) that we had after the binary inspection, but note that we also had in previous works a binary inspection cleaning carried out by a single inspector, resulting in a comparably pure sample. In contrast, we noticed that the 80″80″cutouts help in specific cases to reject non-lenses as the overall environment is visible. Finally, we remark that it is known (Rojas et al., 2023; Schuldt et al., 2025) that every grader has a different expectation and also the grades among a single inspector varies. Consequently, it might be simply that our expectations for an object to qualify as a lens candidate has slightly increased over time, possibly because of the increasing sample of lenses and the higher availability of high-resolution images (e.g., Euclid, see Acevedo Barroso et al. (2024); Nagam et al. (2025), Walmsley et al. in prep.).
4.4 Final lens compilation
Despite the tendency towards lower grades, we adopt our traditional cuts to define grade A () and grade B (), and release in Table 2 the full catalog of visually inspected systems. This catalog includes jointly inspected systems that obtained a committee score and were also detected by VariLens (Andika et al., 2024, noted in column 12 of the catalog), but we exclude any duplicates within ″. Consequently, using the average grades of all eight inspectors to obtain a stable average (see also Rojas et al., 2023; Schuldt et al., 2025), we found 42 grade A and 187 grade B lens candidates, which are listed in Table 2. Furthermore, the newly discovered grade A candidates are shown in Fig. 6, while the grade B systems are shown in Fig. 7. We further provide in this table all grade A and B lens candidates from C21 (marked in column 15, see Table 3), S22 (marked in column 14), and S25 (marked in column 13) which we detected with our network committee as candidate but excluded for visual inspection (see also Sect. 4.2). In sum, we found with our network committee 95 grade A and 503 grade B lenses. From these systems, 506 systems (included not re-graded systems) were already known, while we show in Figs. 8 (grade A) and 9 (grade B) those that we re-graded in our inspection (either because they are in the small comparison sample, see Sect. 4.3, or because they were discovered by other work and not by C21, S22, or S25). This high recovery rate is expected, given the enormous lens search projects that already exploited HSC data and in particular our previous searches using individual networks that entered into our network committee now. However, we remind the reader that the new lens identification is only one aspect of this work.
4.5 Discussion on the lens properties
We further classified the lens environment using photo- histograms during the inspection by the second team (see Fig. 5 for examples), leading to an identification of 147 candidates in an overdensity (corresponding to at least two votes) of which 22 and 66 are ranked as grade A and B candidates, respectively. With the automated photo- procedure to identify overdensities presented by S25, we identify 569 systems among the whole inspected sample, of which 18 and 70 are classified as grade A or B lens candidates, respectively (see also Tab. 1). While the automated procedure can be easily applied to large samples, the visual classification requires in particular human time that is highly limited. On the other hand, we visually identified several systems to be in an overdensity that are missed by the automated procedure due to missing or band observations required for the photo- codes. In total, we find a good overlap between the two complementary procedures.
As mentioned earlier, we further showed the network predicted lens center and Einstein radius during final inspection. Since we found in the calibration round that the model network from Schuldt et al. (2023a) has difficulties to correctly predict the model if the lens is not well centered, we mitigated this through a re-centering between the last two inspection stages. We overlay the predicted lens center and Einstein radius on the cutouts of our lens candidates in Appendix B. For the reported lens center coordinates in Table 2, we only found 9% of the lens candidates to have a relatively poor predicted lens center or Einstein radius which might be because of a group- or cluster-scale lens not suited for the modeling network or a remaining offset of the lens. We provide the predicted values of the seven mass model parameters with their corresponding 1- uncertainties in Table 2 (columns 16 to 29, see also Table 3) as well as a flag (column 30) if the lens center and Einstein radius are matching reasonably the system (defined as two or more votes). This demonstrates once more the power of the modeling network and shows that the result can be used for further analysis.
While we applied in Sect. 3.2 relatively weak cuts on the model parameter uncertainties (see Fig. 4), we note that more stringent cuts are possible when the sample sizes are becoming significantly larger. This would lead to an even higher fraction of systems that can be easily ruled out, while keeping most of the lens candidates.
5 Summary and Conclusion
We have carried out a systematic search for strong gravitational lenses using the bands of the HSC Wide layer observations from the third public data release. For this, we tested the combination of different networks by averaging or multiplying their individual network scores. The best performance, 75% completeness on real lenses at a FPR of , was obtained by averaging the scores from five different networks.
While the resulting sample size would have been easily possible to visually inspect, this will change with the next generation of wide-field imaging surveys, such that further cleaning is unavoidable. We tested two approaches, first using SExtractor and the HSC pixel-level flags to reject images with mostly non-zero background or image artefacts, and second, the ResNet modeling network from Schuldt et al. (2023a) as it predicts higher uncertainties for non-lenses. This lowered the contaminants by 29%, while not excluding any lens candidate from our previous identifications. Such post-processing scripts are expected to play a crucial role when two orders of magnitude more images need to be classified.
Thanks to a visual inspection of the cleaned network candidate list, we identified 95 grade A and 503 grade B candidates, which include 506 previously known systems. We provide in Table 2 their coordinates together with their SIE and external shear parameters obtained from the ResNet presented by Schuldt et al. (2023a). During visual inspection, we also showed, which was done for the first time, the predicted lens center and Einstein radius and characterized the reliability. We found that only % of the candidates obtained a lens center or Einstein radius that is not well predicted, once the system is well centered. Moreover, we characterized their environment with two complementary approaches, either through visual inspection of histograms showing the photo- values of their surrounding, or the selection cuts elaborated by S25. In both cases, we find 88 grade A or B systems to be in a significant overdense environment. By releasing Table 2, which includes all visually inspected candidates with our obtained average grades and their characteristics as well as their previous discoveries, we release one of the most complete catalogs of strong lensing systems so far.
Particularly in preparation for the ongoing and upcoming wide-field imaging surveys, the removal of false-positives of network classifiers will be a challenge, such that the approaches presented here are expected to play a crucial role. Furthermore, fast and autonomous techniques to characterize and analyze the new systems, as tested here as well, will be the challenge we face.
Data Availability
Table 2 will be available after acceptance of the publication at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. It will be further released through the HOLISMOKES collaboration webpage at www.holismokes.org and accordingly incorporated in the SuGOHI data base accessible through https://www-utap.phys.s.u-tokyo.ac.jp/~oguri/sugohi/.
Acknowledgements.
SS has received funding from the European Union’s Horizon 2022 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101105167 — FASTIDIoUS. We acknowledge financial support through grants PRIN-MIUR 2017WSCC32 and 2020SKSTHZ. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (LENSNOVA: grant agreement No 771776). This research is supported in part by the Excellence Cluster ORIGINS which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094 – 390783311. This work uses the following software packages: Astropy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018), matplotlib (Hunter, 2007), NumPy (van der Walt et al., 2011; Harris et al., 2020), Python (Van Rossum & Drake, 2009), Scipy (Virtanen et al., 2020).References
- Acebron et al. (2022) Acebron, A., Grillo, C., Bergamini, P., et al. 2022, A&A, 668, A142
- Acebron et al. (2024) Acebron, A., Grillo, C., Suyu, S. H., et al. 2024, ApJ, 976, 110
- Acevedo Barroso et al. (2024) Acevedo Barroso, J. A., O’Riordan, C. M., Clément, B., et al. 2024, arXiv e-prints, arXiv:2408.06217
- Andika et al. (2024) Andika, I. T., Schuldt, S., Suyu, S. H., et al. 2024, arXiv e-prints, arXiv:2412.12709
- Andika et al. (2023) Andika, I. T., Suyu, S. H., Cañameras, R., et al. 2023, A&A, 678, A103
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
- Bertin & Arnouts (1996) Bertin, E. & Arnouts, S. 1996, A&AS, 117, 393
- Bolton et al. (2004) Bolton, A. S., Burles, S., Schlegel, D. J., Eisenstein, D. J., & Brinkmann, J. 2004, AJ, 127, 1860
- Bolton et al. (2008) Bolton, A. S., Treu, T., Koopmans, L. V. E., et al. 2008, ApJ, 684, 248
- Bonvin et al. (2017) Bonvin, V., Courbin, F., Suyu, S. H., et al. 2017, MNRAS, 465, 4914
- Cañameras et al. (2024) Cañameras, R., Schuldt, S., Shu, Y., et al. 2024, A&A, 692, A72
- Cañameras et al. (2021) Cañameras, R., Schuldt, S., Shu, Y., et al. 2021, A&A, 653, L6
- Cañameras et al. (2020) Cañameras, R., Schuldt, S., Suyu, S. H., et al. 2020, A&A, 644, A163
- Cabanac et al. (2007) Cabanac, R. A., Alard, C., Dantel-Fort, M., et al. 2007, A&A, 461, 813
- Caminha et al. (2019) Caminha, G. B., Rosati, P., Grillo, C., et al. 2019, A&A, 632, A36
- Cao et al. (2020) Cao, X., Li, R., Shu, Y., et al. 2020, MNRAS, 499, 3610
- Chan et al. (2020) Chan, J. H. H., Suyu, S. H., Sonnenfeld, A., et al. 2020, A&A, 636, A87
- Courbin et al. (2018) Courbin, F., Bonvin, V., Buckley-Geer, E., et al. 2018, A&A, 609, A71
- Diehl et al. (2017) Diehl, H. T., Buckley-Geer, E. J., Lindgren, K. A., et al. 2017, ApJS, 232, 15
- Enzi et al. (2024) Enzi, W. J. R., Krawczyk, C. M., Ballard, D. J., & Collett, T. E. 2024, arXiv e-prints, arXiv:2411.08565
- Ertl et al. (2024) Ertl, S., Schuldt, S., Suyu, S. H., et al. 2024, A&A, 685, A15
- Gavazzi et al. (2014) Gavazzi, R., Marshall, P. J., Treu, T., & Sonnenfeld, A. 2014, ApJ, 785, 144
- Gong et al. (2019) Gong, Y., Liu, X., Cao, Y., et al. 2019, ApJ, 883, 203
- Grespan et al. (2024) Grespan, M., Thuruthipilly, H., Pollo, A., et al. 2024, A&A, 688, A34
- Grillo et al. (2024) Grillo, C., Pagano, L., Rosati, P., & Suyu, S. H. 2024, A&A, 684, L23
- Grillo et al. (2018) Grillo, C., Rosati, P., Suyu, S. H., et al. 2018, ApJ, 860, 94
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357–362
- He et al. (2015) He, K., Zhang, X., Ren, S., & Sun, J. 2015, Deep Residual Learning for Image Recognition
- Holloway et al. (2024) Holloway, P., Marshall, P. J., Verma, A., et al. 2024, MNRAS, 530, 1297
- Holwerda et al. (2015) Holwerda, B. W., Baldry, I. K., Alpaslan, M., et al. 2015, MNRAS, 449, 4277
- Hsieh & Yee (2014) Hsieh, B. C. & Yee, H. K. C. 2014, ApJ, 792, 102
- Huang et al. (2021) Huang, X., Storfer, C., Gu, A., et al. 2021, ApJ, 909, 27
- Huang et al. (2020) Huang, X., Storfer, C., Ravi, V., et al. 2020, ApJ, 894, 78
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Inami et al. (2017) Inami, H., Bacon, R., Brinchmann, J., et al. 2017, A&A, 608, A2
- Ivezic et al. (2008) Ivezic, Z., Axelrod, T., Brandt, W. N., et al. 2008, Serbian Astronomical Journal, 176, 1
- Ivezić et al. (2019) Ivezić, Ž., Kahn, S. M., Tyson, J. A., et al. 2019, ApJ, 873, 111
- Jacobs et al. (2019) Jacobs, C., Collett, T., Glazebrook, K., et al. 2019, ApJS, 243, 17
- Jacobs et al. (2017) Jacobs, C., Glazebrook, K., Collett, T., More, A., & McCarthy, C. 2017, MNRAS, 471, 167
- Jaelani et al. (2020) Jaelani, A. T., More, A., Sonnenfeld, A., et al. 2020, MNRAS, 494, 3156
- Jaelani et al. (2024) Jaelani, A. T., More, A., Wong, K. C., et al. 2024, MNRAS, 535, 1625
- Kelly et al. (2023) Kelly, P. L., Rodney, S., Treu, T., et al. 2023, Science, 380, abh1322
- Lange et al. (2024) Lange, S. C., Amvrosiadis, A., Nightingale, J. W., et al. 2024, arXiv e-prints, arXiv:2410.12987
- Lanusse et al. (2018) Lanusse, F., Ma, Q., Li, N., et al. 2018, MNRAS, 473, 3895
- Laureijs et al. (2011) Laureijs, R., Amiaux, J., Arduini, S., et al. 2011, arXiv e-prints, arXiv:1110.3193
- Li et al. (2021) Li, R., Napolitano, N. R., Spiniello, C., et al. 2021, ApJ, 923, 16
- Li et al. (2020) Li, R., Napolitano, N. R., Tortora, C., et al. 2020, ApJ, 899, 30
- Li et al. (2019) Li, R., Shu, Y., Su, J., et al. 2019, MNRAS, 482, 313
- Meštrić et al. (2022) Meštrić, U., Vanzella, E., Zanella, A., et al. 2022, MNRAS, 516, 3532
- Millon et al. (2020) Millon, M., Courbin, F., Bonvin, V., et al. 2020, A&A, 640, A105
- More et al. (2016) More, A., Verma, A., Marshall, P. J., et al. 2016, MNRAS, 455, 1191
- Morishita et al. (2024) Morishita, T., Stiavelli, M., Grillo, C., et al. 2024, ApJ, 971, 43
- Nagam et al. (2025) Nagam, B. C., Acevedo Barroso, J. A., Wilde, J., et al. 2025, arXiv e-prints, arXiv:2502.09802
- Paraficz et al. (2016) Paraficz, D., Courbin, F., Tramacere, A., et al. 2016, A&A, 592, A75
- Pascale et al. (2025) Pascale, M., Frye, B. L., Pierel, J. D. R., et al. 2025, ApJ, 979, 13
- Petrillo et al. (2019) Petrillo, C. E., Tortora, C., Vernardos, G., et al. 2019, MNRAS, 484, 3879
- Planck Collaboration et al. (2020) Planck Collaboration, Aghanim, N., Akrami, Y., et al. 2020, A&A, 641, A6
- Price-Whelan et al. (2018) Price-Whelan, A. M., Sipőcz, B. M., Günther, H. M., et al. 2018, AJ, 156, 123
- Refsdal (1964) Refsdal, S. 1964, MNRAS, 128, 307
- Rojas et al. (2023) Rojas, K., Collett, T. E., Ballard, D., et al. 2023, MNRAS, 523, 4413
- Rojas et al. (2022) Rojas, K., Savary, E., Clément, B., et al. 2022, A&A, 668, A73
- Savary et al. (2022) Savary, E., Rojas, K., Maus, M., et al. 2022, A&A, 666, A1
- Schuldt et al. (2025) Schuldt, S., Cañameras, R., Andika, I. T., et al. 2025, A&A, 693, A291
- Schuldt et al. (2023a) Schuldt, S., Cañameras, R., Shu, Y., et al. 2023a, A&A, 671, A147
- Schuldt et al. (2019) Schuldt, S., Chirivì, G., Suyu, S. H., et al. 2019, A&A, 631, A40
- Schuldt et al. (2023b) Schuldt, S., Suyu, S. H., Cañameras, R., et al. 2023b, A&A, 673, A33
- Schuldt et al. (2021) Schuldt, S., Suyu, S. H., Cañameras, R., et al. 2021, A&A, 651, A55
- Shajib et al. (2021) Shajib, A. J., Treu, T., Birrer, S., & Sonnenfeld, A. 2021, MNRAS, 503, 2380
- Shajib et al. (2022) Shajib, A. J., Wong, K. C., Birrer, S., et al. 2022, A&A, 667, A123
- Shu et al. (2016) Shu, Y., Bolton, A. S., Kochanek, C. S., et al. 2016, ApJ, 824, 86
- Shu et al. (2017) Shu, Y., Brownstein, J. R., Bolton, A. S., et al. 2017, ApJ, 851, 48
- Shu et al. (2022) Shu, Y., Cañameras, R., Schuldt, S., et al. 2022, A&A, 662, A4
- Shu et al. (2018) Shu, Y., Marques-Chaves, R., Evans, N. W., & Pérez-Fournon, I. 2018, MNRAS, 481, L136
- Sonnenfeld et al. (2018) Sonnenfeld, A., Chan, J. H. H., Shu, Y., et al. 2018, PASJ, 70, S29
- Sonnenfeld et al. (2020) Sonnenfeld, A., Verma, A., More, A., et al. 2020, A&A, 642, A148
- Spergel et al. (2015) Spergel, D., Gehrels, N., Baltay, C., et al. 2015, arXiv e-prints, arXiv:1503.03757
- Stein et al. (2022) Stein, G., Blaum, J., Harrington, P., Medan, T., & Lukić, Z. 2022, ApJ, 932, 107
- Stiavelli et al. (2023) Stiavelli, M., Morishita, T., Chiaberge, M., et al. 2023, ApJ, 957, L18
- Storfer et al. (2024) Storfer, C., Huang, X., Gu, A., et al. 2024, ApJS, 274, 16
- Suyu & Halkola (2010) Suyu, S. H. & Halkola, A. 2010, A&A, 524, A94
- Suyu et al. (2012a) Suyu, S. H., Hensel, S. W., McKean, J. P., et al. 2012a, ApJ, 750, 10
- Suyu et al. (2020) Suyu, S. H., Huber, S., Cañameras, R., et al. 2020, A&A, 644, A162
- Suyu et al. (2012b) Suyu, S. H., Treu, T., Blandford, R. D., et al. 2012b, arXiv e-prints, arXiv:1202.4459
- Talbot et al. (2021) Talbot, M. S., Brownstein, J. R., Dawson, K. S., Kneib, J.-P., & Bautista, J. 2021, MNRAS, 502, 4617
- Tanaka et al. (2018) Tanaka, M., Coupon, J., Hsieh, B.-C., et al. 2018, PASJ, 70, S9
- van der Walt et al. (2011) van der Walt, S., Colbert, S. C., & Varoquaux, G. 2011, Computing in Science Engineering, 13, 22
- Van Rossum & Drake (2009) Van Rossum, G. & Drake, F. L. 2009, Python 3 Reference Manual (Scotts Valley, CA: CreateSpace)
- Vanzella et al. (2021) Vanzella, E., Caminha, G. B., Rosati, P., et al. 2021, A&A, 646, A57
- Virtanen et al. (2020) Virtanen, P., Gommers, R., Oliphant, T. E., et al. 2020, Nature Methods, 17, 261
- Wang et al. (2022) Wang, H., Cañameras, R., Caminha, G. B., et al. 2022, A&A, 668, A162
- Wong et al. (2022) Wong, K. C., Chan, J. H. H., Chao, D. C. Y., et al. 2022, PASJ, 74, 1209
- Wong et al. (2018) Wong, K. C., Sonnenfeld, A., Chan, J. H. H., et al. 2018, ApJ, 867, 107
- Wong et al. (2020) Wong, K. C., Suyu, S. H., Chen, G. C. F., et al. 2020, MNRAS
- Zhong et al. (2022) Zhong, F., Li, R., & Napolitano, N. R. 2022, Research in Astronomy and Astrophysics, 22, 065014
Appendix A Color-composite images of re-discovered lens candidates
In this section we present the color-composite image stamps of the lens candidates that obtained an average grade above 1.5 (i.e. grade A or B) during our visual inspection, but are already known in the literature. We note that the lens candidates discovered by C21, S22, and S25, which we excluded before visual inspection to lower the amount, are not shown.
Appendix B Color-composite images of lens candidates with their predicted mass model
In this section, we show the color-composite image stamps of the lens candidates in analogy to Figs. 6 and 7, as well as those in Sect. A. Contrary to previous figures, we show here the lens center and Einstein radius for each system predicted by the ResNet of Schuldt et al. (2023a). These models were also shown in one panel during the visual inspection to help the grader classify.